# HACD: Hand-Aware Conditional Diffusion for Monocular Hand-Held Object Reconstruction

### 23 Nov 2023 v1

@@ 3.3 Hand-Constrained Centroid Fixing
Existing works commonly obtain the camera view $\phi_{C}$ and the hand pose $\phi_{H}$ using standard hand pose estimation methods [30, 60]. Without knowing the object pose during inference, we would need to denoise the points in the hand wrist coordinate system. However, directly using a diffusion model to simultaneously learn the object’s shape and pose is very difficult [19]. Thus, we propose a hand-constrained centroid fixing scheme to ease learning.
@@
Centroid-Fixed Diffusion. We first reduce the problem of object pose estimation to centroid prediction, by defining a new object coordinate system. Essentially, while the origin of the object coordinate system is located at the centroid of the object, the axis orientation is shared with the hand wrist frame. Notice that during training, we can directly use the ground-truth object centroid $\mathcal{M}$ to constrain the object point cloud. Therefore, we harness $\mathcal{M}$ to stabilize the diffusion process and make the point cloud fixed at $\mathcal{M}$. In particular, during centroid-fixed diffusion, we re-center the object point cloud to $\mathcal{M}$ and guarantee the added noise to be of zero-mean by
@@
Noteworthy, by keeping the object centroid unchanged, the misalignment error in semantic feature projection due to centroid movements could also be alleviated. Thus, the training behavior becomes more stable as well.
Centroid Prediction for Reverse Diffusion. During the reverse diffusion process, we propose a simple yet effective hand-constrained network $\mathcal{G}$ to predict the translation of the object w.r.t. the hand wrist coordinate system. Given the input RGB image $I$ along with the corresponding hand pose $\phi_{H}$ and camera view $\phi_{C}$, the predicted object centroid is obtained by
@@
As shown in Fig. 2 (II), $\mathcal{G}$ first encodes the hand vertices $P_{H}$ using PointNet [53] to constrain the object centroid in 3D space. Subsequently, the global hand features are combined with the image features extracted from a pre-trained ResNet-18 [32], followed by two parallel multilayer perceptrons (MLPs) to respectively output the 3D and 2D object centroids.
@@ 3.4 Training Objectives
We adopt Point-Voxel CNN (PVCNN) [41] for our diffusion model $\mathcal{D}_{\theta}$, as it has demonstrated great results in 3D object reconstruction tasks [80, 19, 44]. The point cloud along with its corresponding condition $[X_{t},X_{t}^{Z}]\in\mathbb{R}^{(3+C^{\prime})N}$ is thereby fed into $\mathcal{D}_{\theta}$ to predict the noise $\epsilon_{\theta}\in\mathbb{R}^{3N}$, where $C^{\prime}$ denotes the number of hand-aware conditioning channels. Note that partial layers of PVCNN are adapted for dual-stream conditioning and the details are discussed in the supplementary material.
During optimization, we use the MSE loss between model output $\epsilon_{\theta}(X_{t},t,z)$ and noise $\epsilon$:
@@
Notice that the centroid prediction network is separately trained from $\mathcal{D}_{\theta}$. Thereby, the 3D object centroid $\mathcal{M}_{3d}$ in the hand wrist frame and the 2D object centroid $\mathcal{M}_{2d}$ in the normalized device coordinate (NDC) space are both supervised along with a 2D-3D projection loss. Overall, our total loss is defined as
$\begin{split}\mathcal{L}_{centroid}=&~{}\|\mathcal{M}_{3d}-\widehat{\mathcal{M}}_{3d}\|+\lambda_{1}\|\mathcal{M}_{2d}-\widehat{\mathcal{M}}_{2d}\|\\ &~{}+\lambda_{2}\|\mathcal{P}(\widehat{\mathcal{M}}_{3d})-\widehat{\mathcal{M}}_{2d}\|,\end{split}$ (9)
where $\lambda_{1}$ and $\lambda_{2}$ are hyperparameters, $\mathcal{P}$ is the transformation from the hand wrist frame to the NDC space and $\widehat{\bullet}$ refers to predicted results.
@@ 4.1 Datasets and Setup
Datasets. We evaluate our method on three common benchmark datasets, including the synthetic ObMan [30] dataset and two real-world datasets, namely HO3D [29] and MOW [4]. ObMan [30] consists of 2,772 objects from 8 categories taken from ShapeNet [5], with 21K plausible grasps generated using GraspIt [46]. The dataset contains 141K frames for training and 6K frames for testing, which are rendered using Blender on top of random backgrounds. HO3D [29] includes 77,558 frames from 68 sequences, containing 10 different objects from the YCB dataset [3], which are manipulated by 10 different users. The hand and object pose annotations are obtained from multi-camera optimization procedures. MOW [4] consists of 442 images of 121 object templates, with in-the-wild source data collected from EPIC Kitchens [13, 14] and the 100 Days of Hands [62] datasets. We use the same training and testing split as iHOI [73]. Regarding hand pose prediction, we employ Hasson et al. [30] for the ObMan dataset and FrankMocap [60] for HO3D and MOW, in line with iHOI [73].
Evaluation Metrics. Following [73], we report the Chamfer Distance (CD) in mm and F-score at thresholds of 5mm (F-5) and 10mm (F-10) to compare with the state-of-the-art.
4.2 Evaluation on ObMan Dataset
We first present quantitative results on the large-scale synthetic ObMan dataset in Tab. 1. As can be seen, our approach surpasses all state-of-the-art methods by a large margin for F-score and Chamfer Distance. Specifically, compared with the current best-performing method gSDF [9], we achieve a relative improvement of 27.3 % and 13.6 % for the F-5 and F-10 metrics, respectively. Furthermore, we can reduce the CD by 73.5 % compared with iHOI and 57.8 % compared with HO, demonstrating that our reconstructed shape possesses significantly fewer outliers.
We also visualize the reconstructed objects along with the hands in Fig. 3. While SDF-based methods, including iHOI [73] and gSDF [9], tend to result in either over-smoothed and less detailed, or fragmented reconstructions, our approach is able to generate geometrically coherent point clouds with plausible details, even for thin objects and heavily occluded parts.
F-5 $\uparrow$ F-10 $\uparrow$ CD $\downarrow$
@@
Ours 0.56 0.75 0.27
Table 1: Comparison with State of the Arts on ObMan [30]. F-score of 5mm and 10mm, Chamfer Distance (mm) metrics are utilized for evaluation.
HO3D MOW
Finetuning F-5 $\uparrow$ F-10 $\uparrow$ CD $\downarrow$ F-5 $\uparrow$ F-10 $\uparrow$ CD $\downarrow$
@@
Ours 0.36 0.55 0.66 0.25 0.39 1.86
iHOI [73] 0.14 0.27 4.36 0.09* 0.17* 8.43*
Ours 0.22 0.35 1.34 0.22 0.33 3.15
@@
4.3 Evaluation on Real-World Datasets
Aside from the synthetic ObMan dataset, we also conduct experiments on the two real-world datasets HO3D [29] and MOW [4] to demonstrate our approach’s generalization capabilities for real-world scenarios. The model is respectively finetuned on the two datasets, starting from the ObMan pre-trained model as initialization. As shown in Tab. 2 (top), our approach achieves state-of-the-art results on HO3D as well as MOW. Concretely, regarding F-5 and F-10 metrics, we exhibit a noticeable average improvement of 19.3 % on the HO3D dataset and 77.4 % on the MOW dataset. Compared with the F-score metric, the Chamfer Distance metric is known to be more vulnerable to outliers [65, 73, 9]. Thus, the significant reduction of the CD metric illustrates that our approach is capable of producing significantly fewer outliers. The qualitative results in Fig. 4 also demonstrate the superiority of our approach.
Zero-shot transfer to HO3D and MOW. To further evaluate our zero-shot transfer abilities, we also directly apply our model, trained on the ObMan dataset, to HO3D and MOW without conducting any additional finetuning. The results in Tab. 2 (bottom) show that our method can again achieve a remarkable improvement on HO3D (57 % and 30 % in F-5 and F-10) and MOW (144 % and 94 % in F-5 and F-10), demonstrating the strength of HACD in synthetic-to-real generalization.
4.4 Ablation Studies

### 18 Mar 2024 v2

@@ 3.2 Hand-Constrained Centroid-Fixed Conditional Diffusion
Existing works [78, 80] commonly obtain the camera view $\phi_{C}$ and the hand pose $\phi_{H}$ from standard hand pose estimation methods [31, 61]. Without knowing the object pose during inference, we can only attempt to denoise the points in the hand wrist coordinate system. However, directly using a diffusion model to jointly learn the object’s shape and pose can become very challenging [16]. Thus, we propose a hand-constrained centroid fixing scheme to ease learning.
@@
Centroid-Fixed Diffusion. We first reduce the problem of object pose estimation to centroid prediction, which defines a new object coordinate system. Essentially, while the origin of the object coordinate system is located at the centroid of the object, the axis orientation is shared with the hand wrist frame. Notice that during training, we can directly use the ground-truth object centroid $\mathcal{M}$ to constrain the object point cloud. Therefore, we harness $\mathcal{M}$ to stabilize the diffusion process and make the point cloud fixed at $\mathcal{M}$. In particular, during the centroid-fixed diffusion, we re-center the object point cloud to $\mathcal{M}$ and guarantee that the added noise has zero-mean via
@@
Noteworthy, by keeping the object centroid unchanged, the misalignment error within the semantic feature projection due to centroid movements could also be alleviated. Thus, the training behavior becomes more stable as well.
Centroid Prediction for Reverse Diffusion. During the reverse diffusion process, we propose the use of a simple yet effective network $\mathcal{G}$ to estimate the translation of the object w.r.t. the hand wrist coordinate system. Given the input RGB image $I$ along with the corresponding hand pose $\phi_{H}$ and camera view $\phi_{C}$, the predicted object centroid is obtained as
@@
As shown in Fig. 2 (II), $\mathcal{G}$ first encodes the hand vertices $X_{H}$ using a PointNet-like [54] model to constrain the object centroid in 3D space. Subsequently, the global hand features are combined with the image features, extracted by a pre-trained ResNet-18 [32], and processed by two parallel Multilayer Perceptrons (MLPs) to respectively output the 3D and 2D object centroid.
@@
Noteworthy, estimating object orientation is more challenging than predicting only the translation and will significantly increase the complexity of the network. Therefore, we only constrain the object centroid since previous works [44, 41] have shown diffusion models are able to reconstruct the object well even if it is not in a canonical orientation.
3.3 Dual-Stream Conditional Point Cloud Denoising
The essence of the conditional point cloud denoiser $\mathcal{D}_{\theta}(X_{t},t,z)$ is the modeling of the condition $z$ so to fully utilize the information of the input image along with the corresponding camera view and hand pose. Our key insight is that the hand can offer both semantic and geometric priors to facilitate object reconstruction. Therefore, we utilize a dual-stream architecture to individually process these priors to avoid mutual interference.
Unified Hand-Object Semantic Embedding. Related work has shown that using 2D image features is a vital component for object reconstruction [74, 75]. Unlike prior works that extract a global embedding for the image [12, 74, 75], we instead extract unique deep image features for each point of the partially denoised point cloud at each diffusion step in a projective manner. Specifically, we first extract image features $\mathcal{F}\in\mathbb{R}^{H\times W\times C}$ using a standard 2D network, such as ResNet [32] or ViT [18], with $C$ being the number of output feature channels. We then project the point cloud onto the image by the efficient point cloud rasterizer [57] $\mathcal{R}(\phi_{C},X_{t})$. Subsequently, the pointwise semantic features at timestep $t$ can be represented as $X_{t}^{O}=\pi(\mathcal{R}(\phi_{C},X_{t}),\mathcal{F})$ with $X_{t}^{O}\in\mathbb{R}^{N\times C}$, where $\pi(\cdot)$ denotes the back-projection to 3D space. In this way, every individual point is pixel-aligned with the deep features corresponding to the pixel onto which the point is rasterized.
However, due to the inevitable occlusion induced by the hand, it is very challenging to reconstruct the occluded part of the object as only a single view is provided. On the other side, the hand information should not be fully ignored as the hand naturally serves as a prior for estimating the shape of the hand-occluded object region. Therefore, we apply $\pi(\mathcal{R(\cdot)})$ to both object and hand points to obtain $X_{t}^{HO}=\pi(\mathcal{R}(\phi_{C},[X_{t},X_{H}]),\mathcal{F})$ with $X_{t}^{HO}\in\mathbb{R}^{(N+N_{H})\times C}$, serving as a unified hand-object semantic embedding. Thereby, $N_{H}$ denotes the number of the hand mesh vertices $X_{H}$, which can be obtained from the estimated hand pose $\phi_{H}$ together with the MANO hand model [60]. Noteworthy, we also apply an extra one-hot encoding to indicate whether the points belong to the object or the hand, making $X_{t}^{HO}\in\mathbb{R}^{(N+N_{H})\times(C+1)}$.
Compared with existing works [78, 80], our proposed unified hand-object semantic embedding $X_{t}^{HO}$ holds information of the hand-induced occlusion, which is crucial for the reconstruction of the occluded part of the object, thus increasing the robustness (see Tab. 1 (b)).
Hand Articulation Geometric Embedding. Furthermore, the object shape is also highly constrained by the hand and very geometrically related to the hand articulation. Inspired by [78, 80, 8], we explicitly encode hand-object interaction by transforming the partially denoised point cloud at every step to each hand joint frame.
Given the partially denoised point cloud $X_{t}$ and hand parameters $\phi_{H}$, we first compute the rotation $R_{j}$ and translation $T_{j}$ of each joint $j$ with respective to the hand wrist using forward kinematics given the hand model. We then transform $X_{t}$ from the hand wrist coordinate system to each hand joint frame via $X_{t}^{j}=R_{j}X_{t}+T_{j}\in\mathbb{R}^{N\times 3}$ to encode the hand articulation onto object points. Finally, the hand articulation embedding $X_{t}^{A}\in\mathbb{R}^{N\times J}$ is calculated via concatenation and flattening of $X_{t}^{j}$, with $J$ equaling 45 in our experiments as 15 hand joints are utilized.
Dual-Stream Denoiser. The naive way to utilize the aforementioned semantic and geometric priors is to directly concatenate them as the condition $z=[X_{t}^{HO},X_{t}^{A}]$ and produce the denoised point cloud controlled by $z$ in a single-stream manner. However, forcibly integrating these embeddings from different domains may result in interference and cause a drop in performance (see Tab. 4).
Hence, we instead employ a dual-stream denoiser to separately process $X_{t}^{HO}$ and $X_{t}^{A}$. As shown in Fig. 2, given $z_{1}=X_{t}^{HO}$ and $z_{2}=X_{t}^{A}$, we first feed the object and hand points along with their corresponding semantic embedding $[X_{t},z_{1}]$ to one branch of the dual-stream denoiser $f_{\theta}^{1}$ and obtain $\mathcal{F}_{\theta}^{1}\in\mathbb{R}^{(N+N_{H})\times S}$ as feature representation guided by the semantic prior, with $S$ denoting the number of latent feature channels. Similarly, we feed the object point cloud along with its corresponding geometric embedding $[X_{t},z_{2}]$ to $f_{\theta}^{2}$ using the identical architecture to obtain $\mathcal{F}_{\theta}^{2}\in\mathbb{R}^{N\times S}$ as our geometric feature. The final noise is then predicted from the concatenation of the semantic and geometric features with $\epsilon_{\theta}=g_{\theta}([\mathcal{F}_{\theta}^{1},\mathcal{F}_{\theta}^{2}])\in\mathbb{R}^{N\times 3}$, where $g_{\theta}$ consists of stacked Multilayer Perceptrons (MLPs).
In this way, each branch can learn a specialized representation suitable for its own domain and then serve as conditioning to the diffusion model, contributing to the reconstruction of object shape from different domains. The detailed architecture of the dual-stream denoiser is provided in supplementary material.
@@ 3.4 Training Objectives
Diffusion model. For optimization, we use the common MSE loss between model prediction $\epsilon_{\theta}(X_{t},t,z)$ and applied noise $\epsilon$ with
@@
We further regularize the object shape using a projective mask loss $\mathcal{L}_{mask}$. To this end, at each timestep $t$ the sampled and predicted point cloud are projected onto the image via the aforementioned rasterizer, and the L1 loss between them is computed as
$\mathcal{L}_{mask}=\|\mathcal{R}(X_{t})-\mathcal{R}(\hat{X_{t}})\|_{1},$ (9)
where $\hat{\bullet}$ denotes predicted results. The overall loss is then a weighted sum of both terms with
$\mathcal{L}_{overall}=\mathcal{L}_{denoise}+\eta_{1}\mathcal{L}_{mask},$ (10)
where $\eta_{1}$ is a hyperparameter that controls the strength of the projective 2D regularization term.
Centroid prediction network. Notice that the centroid prediction network is separately trained from $\mathcal{D}_{\theta}$. Thereby, the 3D object centroid $\mathcal{M}_{3d}$ in the hand wrist frame and the 2D object centroid $\mathcal{M}_{2d}$ in the normalized device coordinate (NDC) space are both supervised along with a 2D-3D projection loss. Overall, our total loss is defined as
$\begin{split}\mathcal{L}_{centroid}=\|\mathcal{M}_{3d}-\widehat{\mathcal{M}}_{3d}\|+\lambda_{1}\|\mathcal{M}_{2d}-\widehat{\mathcal{M}}_{2d}\|+\lambda_{2}\|\mathcal{P}(\widehat{\mathcal{M}}_{3d})-\widehat{\mathcal{M}}_{2d}\|,\end{split}$ (11)
where $\lambda_{1}$ and $\lambda_{2}$ are hyperparameters, $\mathcal{P}$ is the transformation from the hand wrist frame to the NDC space, and $\widehat{\bullet}$ refers to predicted results.
@@ 4.1 Datasets and Setup
Datasets. We evaluate our method on four common benchmark datasets, including the synthetic ObMan [31] dataset and three real-world datasets, namely HO3D [28], MOW [4] and DexYCB [6]. ObMan [31] consists of 2,772 objects from 8 categories taken from ShapeNet [5], with 21K plausible grasps generated using GraspIt [46]. The dataset contains 141K frames for training and 6K frames for testing, which are rendered using Blender on top of random backgrounds. HO3D [28] includes 77,558 frames from 68 sequences, containing 10 different objects from the YCB dataset [3], which are manipulated by 10 different users. The hand and object pose annotations are obtained from multi-camera optimization procedures. MOW [4] consists of 442 images of 121 object templates, with in-the-wild source data collected from EPIC Kitchens [15, 14] and the 100 Days of Hands [65] datasets. We use the same training and testing split as iHOI [78] and DDF-HO [80] for all three aforementioned datasets. DexYCB [6] is a large-scale real-world hand-object dataset. Following [9, 76, 8], we focus on right-hand samples and use the official s0 split. 29,656 training samples and 5,928 testing samples are downsampled following the setup of gSDF[8].
Table 1: Comparison with the state-of-the-art on ObMan [31]. (a) F-score of 5mm and 10mm, Chamfer Distance (mm) metrics are utilized for evaluation. (b) Robustness against hand occlusion. We analyze the patterns of the F-5 and F-10 metrics as a function of the object visibility ratio on the test set of ObMan.
Method F-5 $\uparrow$ F-10 $\uparrow$ CD $\downarrow$
@@
DDF-HO [80] 0.55 0.67 0.14
Ours 0.61 0.81 0.11
Evaluation Metrics. Following [78, 80], we report the Chamfer Distance (CD) in mm and F-score at thresholds of 5mm (F-5) and 10mm (F-10) to compare with the state-of-the-art.
Implementation Details are provided in supplementary material.
4.2 Evaluation on ObMan Dataset
We first present quantitative results on the large-scale synthetic ObMan dataset in Tab. 1 (a). As can be seen, our approach surpasses all state-of-the-art methods by a large margin for F-score and Chamfer Distance. Specifically, compared with the current best-performing method DDF-HO [80], we achieve a relative improvement of 10.9 % and 20.9 % for the F-5 and F-10 metrics, respectively. Furthermore, we can reduce the CD by 21.4 % compared with DDF-HO and 89.2 % compared with iHOI, demonstrating that our reconstructed shape possesses significantly fewer outliers.
We also visualize the reconstructed objects along with the hands in Fig. 3. Since we focus on object reconstruction, we use ground-truth hand poses for all images. The object surfaces are reconstructed using alpha shapes [19] implemented with Open3D [84] to qualitatively compare with previous works. Noteworthy, the metrics are not affected as they are computed from the point clouds. While SDF- or DDF-based methods, including iHOI [78], gSDF [8] and DDF-HO [80], tend to result in either over-smoothed and less-detailed, or fragmented reconstructions, our approach is able to generate geometrically coherent point clouds with plausible details, even for thin objects and heavily occluded parts. Specifically, the bottle in Row 2 suffers a 54.4% occlusion rate. Nonetheless, D-SCo shows strong capabilities of inferring the occluded part of the object.
Method Finetuning HO3D MOW
@@ F-5 $\uparrow$ F-10 $\uparrow$ CD $\downarrow$ F-5 $\uparrow$ F-10 $\uparrow$ CD $\downarrow$
DDF-HO [80]  [^1] The results are provided by the authors of DDF-HO [80], with exact same training split as iHOI [78] and ours. 0.27 0.40 0.86 0.17 0.24 1.59
Ours 0.41 0.63 0.34 0.30 0.47 1.35
iHOI [78] 0.14 0.27 4.36 0.09 0.17 8.43
DDF-HO [80] 0.24 0.36 0.73 0.14 0.19 1.89
Ours 0.27 0.41 1.19 0.24 0.36 1.44
4.3 Evaluation on Real-World Datasets
Aside from the synthetic ObMan dataset, we also conduct experiments on the three real-world datasets HO3D [28], MOW [4] and DexYCB [6] to demonstrate our approach’s generalization capabilities for real-world scenarios. The model is respectively finetuned on the HO3D and MOW, starting from the ObMan pre-trained model as initialization. As shown in Tab. 2 (top) and Tab. 3, our approach achieves state-of-the-art results on HO3D, MOW as well as DexYCB. Concretely, regarding F-5 and F-10 metrics, we exhibit a noticeable average improvement of 54.7 %, 86.2 %, and 29.4 %, respectively. In contrast to the F-score metric, the Chamfer Distance is known to be more vulnerable to outliers [68, 78, 8]. Thus, the significant reduction of the CD metric illustrates that our approach is significantly more robust, producing fewer outliers. The qualitative results in Fig. 4 similarly demonstrate the superiority of our approach.
Zero-shot transfer to HO3D and MOW. To further evaluate our zero-shot transfer abilities, we also directly apply our model, trained on the ObMan dataset, to HO3D and MOW without conducting any additional finetuning. The results in Tab. 2 (bottom) show that our method can again achieve a remarkable improvement on HO3D (13 % and 14 % in F-5 and F-10) and MOW (71 % and 89 % in F-5 and F-10), demonstrating the strength of D-SCo in synthetic-to-real generalization.
Table 3: Comparison with the state-of-the-art on DexYCB [6].
Metric HO [31] GF [37] AlignSDF [9] gSDF [8] Ours
F-5 $\uparrow$ 0.38 0.39 0.41 0.44 0.63
F-10 $\uparrow$ 0.64 0.66 0.68 0.71 0.82
CD $\downarrow$ 0.42 0.45 0.39 0.34 0.13
4.4 Ablation Studies
Effectiveness of dual-stream denoiser. We demonstrate the effectiveness of dual-stream denoiser in Tab. 4. Instead of individually processing semantic and geometric embeddings, we simply concatenate the embeddings as conditioning to the denoiser. The decrease in performance (C0 vs. B0) shows that independently processing semantic and geometric information within our dual-stream denoiser enhances reconstruction quality.
Effectiveness of semantic and geometric condition. In Tab. 4, we show the impact of the semantic and geometric embeddings. The proposed unified hand-object semantic embedding $X_{t}^{HO}$ utilizes the semantic information of the hand to supplement the occluded semantic information of objects, which is crucial to model hand-induced occlusion. Adding the embedding leads again to improved performance (D2 vs. D0, C0 vs. D1). Moreover, the hand articulation geometric embedding $X_{t}^{A}$ explicitly models hand-object interaction geometrically and results in improved performance (D1 vs. D0, C0 vs. D2).
We further noticed that the way of encoding the hand-object conditional information has a significant impact on performance. In particular, we have explored two alternative strategies to model the hand-object interaction. First, inspired by [17], we implement a GCN-based hand embedding, which utilizes a graph convolutional network [22] for feature extraction. We again apply the GCN hand embedding to the partially denoised point cloud at each step. Although the GCN hand embedding implicitly encodes the hand articulation, it does not encode the hand-object interaction. Therefore, simply applying such an embedding as conditioning does not further performance (See D3 vs. C0). The hand articulation geometric embedding, on the other hand, applies the unique articulation-aware embedding to each point of the partially denoised point cloud, which explicitly encodes hand-object interaction, thus significantly benefiting the object reconstruction performance. Second, we also utilize standard PointNet [54] to encode 3D hand vertices into a global hand embedding, which is then served to the partially denoised point cloud. Again, without modeling the hand-object interaction, the global hand embedding falls short of providing sufficient semantic information about the hand-occluded object. Consequently, the predictions end up being inferior to our proposed $X_{t}^{HO}$ (D4 vs. C0).
@@ 4.4 Ablation Studies
Effectiveness of hand-constrained centroid fixing. In Tab. 4, we illustrate the importance of the centroid fixing scheme. Our centroid fixing operation improves the stability of the diffusion and the reverse processes, thus enhancing the performance (D0 vs. E0). Further, without the hand-constrained centroid prediction network, the diffusion model has to simultaneously learn the centroid deviation and object shape, which leads to clearly worse results (E1 vs. D0). Additionally, in F0 we report results when using the actual ground-truth object centroid, and in F1 when using the ground-truth object pose. Without our centroid fixing scheme, the results remain inferior even when using the ground-truth object centroid on ObMan (F0 vs. D0). This further demonstrates the power of our proposed centroid fixing paradigm.
Effectiveness of $\mathcal{L}_{mask}$. We compare the quantitative results w/ or w/o $\mathcal{L}_{mask}$. Supervising the diffusion process in both 2D and 3D domains further boosts the object reconstruction performance (A0 vs. B0). Noteworthy, we can surpass existing methods on both ObMan and HO3D datasets even without $\mathcal{L}_{mask}$ (B0).
Table 4: Ablation study on ObMan [31] and HO3D [28] datasets.